# OODEval: Evaluating Large Language Models on Object-Oriented Design

### v1

@@ 4. CLUE: Class Likeness Unified Evaluation Metric
CLUE is a composite metric set comprising the clue metric for evaluating overall class diagram similarity, clue-class for assessing class similarity, clue-attribute for evaluating class attribute similarity, clue-method for assessing class method similarity, and clue-relation for evaluating class relationship similarity. To prevent ambiguity, uppercase CLUE denotes the entire metric set, whereas lowercase clue refers to a specific individual metric. In comparison to prior approaches, CLUE comprehensively incorporates class diagram structures alongside richer semantic information, while computing finer-grained semantics during the structural matching process, thereby substantially enhancing the performance of the evaluation metrics. The following Definitions 4.7 to 4.10 define the computation methods for these five metrics.
Definition 4.0 (Class Diagram Similarity, clue).
Given a reference model $M_{1}$ and a candidate model $M_{2}$, the class diagram similarity is the weighted average of the class similarity(see Definition 4.8) and the class relationship similarity(see Definition 4.9), which is:
(6) $clue(M_{1},M_{2})=w_{e}\cdot clue\mbox{-}class(M_{1},M_{2})+w_{r}\cdot clue\mbox{-}relation(M_{1},M_{2})$
Definition 4.0 (Class Similarity, clue-class).
Given a reference model $M_{1}$ with $n$ classes and a candidate model $M_{2}$ with $m$ classes, the class similarity $\textit{clue-class}(M_{1},M_{2})$ is defined as $Sim(ES_{ce})$(see Definition 4.2), where $ES_{ce}\in\mathbb{R}^{n\times m}$ is the class similarity matrix constructed as follows:
(7) $ES_{ce}[i][j]=w_{n}\cdot\sigma(c_{1}^{(i)},c_{2}^{(j)})+w_{a}\cdot\text{sim}_{ca}(c_{1}^{(i)},c_{2}^{(j)})+w_{m}\cdot\text{sim}_{cm}(c_{1}^{(i)},c_{2}^{(j)})$
Definition 4.0 (Class Relationship Similarity, clue-relation).
Given a reference model $M_{1}$ with $p$ relationships and a candidate model $M_{2}$ with $q$ relationships, the class relationship similarity $clue\mbox{-}relation(M_{1},M_{2})$ is defined as $Sim(ES_{cr})$, where the class relationship similarity matrix $ES_{cr}\in\mathbb{R}^{p\times q}$ is:
$ES_{cr}[i][j]$ $=w_{rt}\cdot\text{LUT}(r_{1}^{(i)}.\text{r\_type},r_{2}^{(j)}.\text{r\_type})+w_{rq}\cdot\text{sim}_{rq}(r_{1}^{(i)},r_{2}^{(j)})$
$+\frac{w_{rn}}{2}\cdot\left(ES_{ce}[idx(r_{1}^{(i)}.\text{c\_begin})][idx(r_{2}^{(j)}.\text{c\_begin})]\right$.
(8) $\quad\quad\quad\left.+ES_{ce}[idx(r_{1}^{(i)}.\text{c\_end})][idx(r_{2}^{(j)}.\text{c\_end})]\right)$
where $idx(c)$ is the index of class $c$ in $C_{1}$ or $C_{2}$ set. And $\text{LUT}(t_{1},t_{2})\to[0,1]$ is the relationship type similarity, between $t_{1}$ and $t_{2}$ , where $t_{1},t_{2}\in\mathbb{T}$, as defined in (Al-Khiaty and Ahmed, 2014).
@@
Definition 4.0 (Class Attribute/Method Similarity, clue-attribute/clue-method).
Given a reference model $M_{1}$ with $n$ classes and a candidate model $M_{2}$ with $m$ classes, let $\pi:[1\dots\min(n,m)]\to[1\dots\max(n,m)]$ be the optimal class matching derived from the class similarity matrix $ES_{ce}$ (Definition 4.8), where $\pi$ maps indices of the smaller class set to the larger one. The class attribute/method similarity is defined as:
(9) $clue\mbox{-}attribute(M_{1},M_{2})=\frac{1}{n}\sum_{i=1}^{min(n,m)}\text{sim}_{ca}(c_{1}^{(i)},c_{2}^{(\pi(i))})$
(10) $clue\mbox{-}method(M_{1},M_{2})=\frac{1}{n}\sum_{i=1}^{min(n,m)}\text{sim}_{cm}(c_{1}^{(i)},c_{2}^{(\pi(i))})$
@@
Figure 2 illustrates the computation dependency graph for the five aforementioned CLUE metrics. In the figure, green nodes represent the five core metrics of CLUE (clue-class, clue-attribute, clue-method, clue-relation, and the overall clue); blue nodes denote the similarity matrices that require optimal assignment, as formally defined in Equations (3)-(4.9); and gray nodes correspond to the leaf nodes of the recursion, whose values are directly used to populate the entries of the similarity matrices. The computation of CLUE follows a multi-layer recursive process with layer-wise weighting. Starting from the root node clue, the recursion proceeds with a maximum depth of four layers. Each edge in the graph is associated with a tunable weight parameter, resulting in a total of 15 trainable weights across the entire dependency structure.
@@ 4.3. Hyperparameter Optimization
Specifically, $\boldsymbol{x}=(a,b,c,d,e,f,g,h,k)\in\mathbb{R}^{9}$ represents the 9-dimensional decision variable vector parameterizing the weights $w_{e}$, $w_{n}$, $w_{a}$, $w_{at}$, $w_{mp}$, $w_{mn}$, $w_{pt}$, $w_{rt}$, and $w_{rq}$, respectively. The remaining six weights are derived via normalization constraints, partitioning the weights into groups that sum to 1 (e.g., $w_{r}=1-w_{e}$ for pairwise groups and $w_{m}=1-w_{n}-w_{a}$ for three-weight groups), thus avoiding explicit optimization for complementary weights. The feasible region $\mathcal{X}\subseteq\mathbb{R}^{9}$ is defined by the 18-dimensional constraint vector $\boldsymbol{g}(\boldsymbol{x})$, comprising inequalities $g_{1}$ through $g_{18}$ as detailed in Equation 13. The objective function $f(\boldsymbol{x})$ is defined in Equation 12. $\mathcal{M}$ denote the set of evaluation metric types (i.e., clue, clue-attribute, clue-method, clue-class, clue-relation). The mapping $C:\mathbb{R}^{9}\times\mathcal{M}\to\mathbb{R}^{n}$ yields $C_{m}(\boldsymbol{x})$, the CLUE score vector for metric type $m\in\mathcal{M}$ under configuration $\boldsymbol{x}$. The vector $\boldsymbol{y}_{\text{human}}\in\mathbb{R}^{n}$ represents the corresponding human-rated scores on the same samples. The Pearson correlation coefficient $\rho(\boldsymbol{u},\boldsymbol{v}):\mathbb{R}^{n}\times\mathbb{R}^{n}\to[-1,1]$ quantifies the linear correlation between vectors $\boldsymbol{u}$ and $\boldsymbol{v}$.
@@
As shown in Table 5, all metrics demonstrate at least moderate correlations ($\rho>0.3$) with human scores on the training set, while exhibiting robust generalization to the test set. In particular, the clue-class and clue metrics achieve strong correlations ($\rho>0.5$). The p-values for all correlations are substantially below 0.05, indicating that the observed associations are statistically significant and reliable. The resulting optimal weight configuration is: $w_{e}=0.810$, $w_{r}=0.190$, $w_{n}=0.787$, $w_{a}=0.104$, $w_{m}=0.109$, $w_{at}=0.594$, $w_{an}=0.406$, $w_{mn}=0.730$, $w_{mt}=0.153$, $w_{mp}=0.117$, $w_{pt}=0.050$, $w_{pn}=0.950$, $w_{rt}=0.156$, $w_{rq}=0.220$, $w_{rn}=0.624$. This configuration will serve as the baseline for subsequent empirical evaluations.
4.4. Metrics Performance Evaluation
In the performance evaluation of metrics, we selected BLEU(Papineni et al., 2002), CodeBERTScore(Zhou et al., 2023), and NINHS (Al-Khiaty and Ahmed, 2014) as baselines for comparison with CLUE, representing text matching, embedding-based, and domain-specific approaches, respectively. We evaluated the performance of these metrics by calculating correlation coefficients between the baseline methods, CLUE, and human ratings from the OODEval-Human dataset. A higher correlation coefficient indicates closer alignment between the automated evaluation method and human judgments. Table 6 presents the Pearson and Spearman correlation coefficients for the baseline methods, CLUE, and its sub-metrics relative to human ratings, along with their statistical significance. The results demonstrate that: (1) CLUE significantly outperforms BLEU, CodeBERTScore, and NINHS in both Pearson and Spearman correlations with human ratings, achieving highly significant statistical outcomes. Notably, CLUE’s Pearson correlation coefficient exceeds 0.5, reflecting a strong alignment with human judgments and enhanced interpretability(Cohen, 1992). (2) Compared to its sub-metrics (class, attribute, method, relation), the CLUE metric achieves higher correlations, indicating that a comprehensive evaluation integrating names, attributes, methods, and relationships better aligns with human judgments and enhances interpretability compared to individual element-based evaluations.
@@ 4.5. Ablation Study on the Underlying Semantic Model
Since CLUE employs a replaceable semantic model to compute string similarity at the leaf nodes, it is necessary to further examine how different underlying semantic models influence its performance. To this end, we evaluate CLUE across multiple semantic models. Table 7 reports Pearson’s correlation coefficients obtained using four semantic models, namely MiniLM(Wang et al., 2020), BGE(Chen et al., 2023a), BERT(Devlin et al., 2019), and CodeBERT (Feng et al., 2020), on the CLUE metrics as well as on direct embedding similarity. The results show that: (1) CLUE substantially outperforms the direct computation of vector similarity using semantic models, which demonstrates the effectiveness of the CLUE algorithm itself; and (2) the CLUE metrics based on CodeBERT achieve the best overall performance. Although the clue-relation and clue-attribute metrics perform slightly worse than those based on BGE, the differences are minor, with the gap not exceeding 0.025. Therefore, given its superior overall performance, we adopt CodeBERT as the semantic model in the subsequent empirical studies.
Models embedding clue-class clue-attribute clue-method clue-relation clue
MiniLM 0.157629 0.398578 0.377008 0.317816 0.459131 0.435976
BGE 0.249080 0.467757 0.400091* 0.336586 0.462346* 0.496893
BERT 0.332595* 0.573388 0.389749 0.341215 0.460218 0.589596
CodeBERT 0.166168 0.584867* 0.375269 0.388763* 0.447774 0.594927*
* denotes the maximum value
@@ 5. Empirical Study
RQ5: What are the prevalent failure modes of LLMs on OOD tasks?
@@
Model Context Length Local? Open Source? Code LLM? Instruct Fine-Tuned? Reasoning LLM? Publish Time
llama3.1-8b-it 128K 2024-07
llama3.1-70b-it 128K 2024-07
llama2-7b 4K 2023-07
llama2-13b 4K 2023-07
codellama-7b-it 16K 2023-08
codellama-13b-it 16K 2023-08
codellama-34b-it 16K 2023-08
codellama-7b 16K 2023-08
codellama-13b 16K 2023-08
codellama-34b 16K 2023-08
starcoder2-3b 16K 2024-03
starcoder2-7b 16K 2024-03
starcoder2-15b 16K 2024-03
gemma3-4b-it 128K 2025-03
gemma3-12b-it 128K 2025-03
gemma3-27b-it 128K 2025-03
gemma3-4b-pt 128K 2025-03
gemma3-12b-pt 128K 2025-03
gemma3-27b-pt 128K 2025-03
qwen3-8B 32K 2025-04
qwen3-14b 32K 2025-04
qwen3-32b 32K 2025-04
qwen3-coder-30b 256K 2025-07
deepseek-coder-33b 16K 2023-11
deepseek-r1 32K 2025-01
deepseek-v3 128K 2024-07
gpt-3.5-turbo 4K - 2024-03
gpt-4o-mini 128K - 2024-07
gpt-4o 128K - 2024-05

### v2

@@ 4. CLUE: Class Likeness Unified Evaluation Metric
CLUE represents a comprehensive suite of metrics designed to evaluate class diagrams at multiple granularities (see Definition 4.6): the primary clue metric assesses overall diagram similarity; clue-class focuses on individual class similarity; clue-attribute evaluates the similarity of class attributes; clue-method measures the similarity of class methods; and clue-relation quantifies the similarity of inter-class relationships. To avoid nomenclature ambiguity, we use uppercase CLUE to denote the entire metric suite, while lowercase clue specifically refers to the overarching metric for class diagram similarity assessment. Compared to existing methods, CLUE distinctively integrates both structural aspects of class diagrams and enriched semantic representations, while incorporating fine-grained semantic computations within the structural matching phase, thereby enabling a more accurate and fine-grained evaluation.
Definition 4.0 (Class Likeness Unified Evaluation Metric, CLUE).
Given a reference class diagram model $\mathcal{D}_{1}=(C_{1},R_{1})$ and a candidate model $\mathcal{D}_{2}=(C_{2},R_{2})$, where $C_{1}=(c^{1}_{1},\dots,c^{1}_{n})$ and $C_{2}=(c^{2}_{1},\dots,c^{2}_{m})$, each class $c^{1}_{i}\in C_{1}$ is defined as $c^{1}_{i}=(name^{1}_{i},A^{1}_{i},M^{1}_{i})$ and each class $c^{2}_{j}\in C_{2}$ is defined as $c^{2}_{j}=(name^{2}_{j},A^{2}_{j},M^{2}_{j})$. The CLUE metric suite consists of five interrelated similarity metrics defined as:
(5) $clue(\mathcal{D}_{1},\mathcal{D}_{2})=w_{e}\cdot clue\mbox{-}class(\mathcal{D}_{1},\mathcal{D}_{2})+w_{r}\cdot clue\mbox{-}relation(\mathcal{D}_{1},\mathcal{D}_{2})$
(6) $\textit{clue-class}(\mathcal{D}_{1},\mathcal{D}_{2})=OptMatch(C_{1},C_{2})$
(7) $\textit{clue-relation}(\mathcal{D}_{1},\mathcal{D}_{2})=OptMatch(R_{1},R_{2})$
(8) $clue\mbox{-}attribute(\mathcal{D}_{1},\mathcal{D}_{2})=\begin{cases}\frac{1}{n}\sum_{i=1}^{n}OptMatch(A^{1}_{i},A^{2}_{\pi(i)}),&\text{if }m\geq n>0\\ \frac{1}{n}\sum_{j=1}^{m}OptMatch(A^{1}_{\pi^{\prime}(j)},A^{2}_{j}),&\text{if }n>m>0\end{cases}$
(9) $clue\mbox{-}method(\mathcal{D}_{1},\mathcal{D}_{2})=\begin{cases}\frac{1}{n}\sum_{i=1}^{n}OptMatch(M^{1}_{i},M^{2}_{\pi(i)}),&\text{if }m\geq n>0\\ \frac{1}{n}\sum_{j=1}^{m}OptMatch(M^{1}_{\pi^{\prime}(j)},M^{2}_{j}),&\text{if }n>m>0\end{cases}$
where $\pi(\cdot)$ and $\pi^{\prime}(\cdot)$ denote the injective mappings induced by the optimal class matching between $C_{1}$ and $C_{2}$, as defined in Definition 4.4.
@@
Figure 2 illustrates the computation dependency graph for the five aforementioned CLUE metrics. The computation of CLUE follows a multi-layer recursive process with layer-wise weighting. Starting from the root node clue, the recursion proceeds with a maximum depth of four layers. Each recursive edge in the graph is associated with a tunable weight parameter, resulting in a total of 15 trainable weights across the entire dependency structure. In the figure, green nodes represent the five core metrics of CLUE (clue-class, clue-attribute, clue-method, clue-relation, and the overall clue); blue nodes denote the similarity matrices that require optimal matching, as formally defined in Table 5; orange nodes represent the optimal matching similarity on $ES$ matrix between two sets; and gray nodes correspond to the leaf nodes of the recursive computation, each of which implements an atomic similarity function serving as the base case of the CLUE metric.
As illustrated in Figure 2, CLUE requires five types of similarity matrices, namely $ES_{ce}$, $ES_{ca}$, $ES_{cm}$, $ES_{mp}$, and $ES_{cr}$, which are computed over class sets, relationship sets, attribute sets, method sets, and parameter sets, respectively. All these similarity matrices are defined as specific instantiations of the Element Set Similarity Matrix (Definition 4.3) by instantiating the four aspects: (1) the element sets $A$ and $B$; (2) the internal representation $d\mbox{-}tuple$ of each element in $A$ and $B$; (3) the corresponding weight vector $\mathbf{w}$; and (4) the corresponding similarity functions. Table 5 summarizes the instantiation details for these five types of similarity matrices.
Table 5. Instantiation for Element Set Similarity Matrices.
ES Element Sets d-tuple weight vector similarity function
$ES_{ce}$ class sets $(name,A,M)$ $(w_{n},w_{a},w_{m})$ $(\sigma,OptMatch,OptMatch)$
$ES_{ca}$ attribute sets $(name,type)$ $(w_{an},w_{at})$ $(\sigma,\sigma)$
$ES_{cm}$ method sets $(name,returnType,P)$ $(w_{mn},w_{mt},w_{mp})$ $(\sigma,\sigma,OptMatch)$
$ES_{mp}$ parameter sets $(name,type)$ $(w_{pn},w_{pt})$ $(\sigma,\sigma)$
$ES_{cr}$ relationship sets $(type,begin,end,label)$ $(w_{rt},w_{rn}/2,w_{rn}/2,w_{rm})$ $(LUT,ES_{ce},ES_{ce},\operatorname{sim}_{rm})$
all of the notions in the table come from Definition 4.1 to 4.7.
The similarity matrices $ES_{ce}$, $ES_{ca}$, $ES_{cm}$, and $ES_{mp}$ adhere to a unified computational framework. Specifically, the string similarity function $\sigma(\cdot,\cdot)$ is employed for pairwise comparisons of string-valued elements, whereas the optimal matching similarity $\text{OptMatch}(\cdot,\cdot)$ is utilized for comparing sets of elements. The computation proceeds in a bottom-up, recursive pattern: $ES_{mp}$ is calculated first, followed by $ES_{cm}$ and $ES_{ca}$, with the results ultimately aggregated into $ES_{ce}$. This hierarchical design guarantees local optimality at each level, which is then propagated to subsequent higher-level aggregations. In contrast, the relationship similarity matrix $ES_{cr}$ adopts a distinct formulation. This stems from the fact that relationship similarity is contingent not only on intrinsic attributes (e.g., relationship type and multiplicity) but also on the similarity of the classes connected at both endpoints. Consequently, the viability of a relationship match is profoundly influenced by the semantic alignment of its source and target classes. To precisely delineate the associated notation in $ES_{cr}$ and , we formally define the relationship similarity matrix in Definition 4.7.
Definition 4.0 (Relationship Similarity Matrix).
Let $R_{1}=\{r^{1}_{1},\dots,r^{1}_{p}\}$ and $R_{2}=\{r^{2}_{1},\dots,r^{2}_{q}\}$ be the reference and candidate relationship sets, respectively. Each relationship $r=(\textit{type},\textit{begin},\textit{end},\textit{label})$ follows the definition in the class diagram model (Definition 4.1). The relationship similarity matrix is defined as $ES_{cr}\in[0,1]^{p\times q}$, where each entry $ES_{cr}(i,j)$ is computed as:
(10) $ES_{cr}(i,j)=w_{rt}\cdot\mathrm{LUT}(type^{1}_{i},type^{2}_{j})+\frac{w_{rn}}{2}\cdot\left(ES_{ce}(begin^{1}_{i},begin^{2}_{j})+ES_{ce}(end^{1}_{i},end^{2}_{j})\right)+w_{rm}\cdot\mathrm{sim}_{rm}(label^{1}_{i},label^{2}_{j})$
for $i=1,\dots,p$ and $j=1,\dots,q$, where $w_{rt},w_{rm},w_{rn}\geq 0$ and $w_{rt}+w_{rm}+w_{rn}=1$. Here, $\mathrm{LUT}:\mathbb{T}\times\mathbb{T}\rightarrow[0,1]$ denotes the relationship type similarity defined by a lookup table following the configuration in (Al-Khiaty and Ahmed, 2014). The function $\mathrm{sim}_{rm}(\cdot,\cdot)$ measures the similarity between relationship multiplicity labels (Definition 4.5). The term $ES_{ce}(i,j)$ refers to the similarity between the $i$-th reference class and the $j$-th candidate class in the class similarity matrix.
@@ 4.3. Hyperparameter Optimization
Specifically, $\boldsymbol{x}=(a,b,c,d,e,f,g,h,k)\in\mathbb{R}^{9}$ represents the 9-dimensional decision variable vector parameterizing the weights $w_{e}$, $w_{n}$, $w_{a}$, $w_{at}$, $w_{mp}$, $w_{mn}$, $w_{pt}$, $w_{rt}$, and $w_{rm}$, respectively. The remaining six weights are derived via normalization constraints, partitioning the weights into groups that sum to 1 (e.g., $w_{r}=1-w_{e}$ for pairwise groups and $w_{m}=1-w_{n}-w_{a}$ for three-weight groups), thus avoiding explicit optimization for complementary weights. The feasible region $\mathcal{X}\subseteq\mathbb{R}^{9}$ is defined by the 18-dimensional constraint vector $\boldsymbol{g}(\boldsymbol{x})$, comprising inequalities $g_{1}$ through $g_{18}$ as detailed in Equation 13. The objective function $f(\boldsymbol{x})$ is defined in Equation 12. $\mathcal{M}$ denote the set of evaluation metric types (i.e., clue, clue-attribute, clue-method, clue-class, clue-relation). The mapping $C:\mathbb{R}^{9}\times\mathcal{M}\to\mathbb{R}^{n}$ yields $C_{m}(\boldsymbol{x})$, the CLUE score vector for metric type $m\in\mathcal{M}$ under configuration $\boldsymbol{x}$. The vector $\boldsymbol{y}_{\text{human}}\in\mathbb{R}^{n}$ represents the corresponding human-rated scores on the same samples. The Pearson correlation coefficient $\rho(\boldsymbol{u},\boldsymbol{v}):\mathbb{R}^{n}\times\mathbb{R}^{n}\to[-1,1]$ quantifies the linear correlation between vectors $\boldsymbol{u}$ and $\boldsymbol{v}$.
@@
Models clue-class clue-attribute clue-method clue-relation clue
MiniLM 0.398578 0.377008 0.317816 0.459131 0.435976
BGE 0.467757 0.400091* 0.336586 0.462346* 0.496893
BERT 0.573388 0.389749 0.341215 0.460218 0.589596
CodeBERT 0.584867* 0.375269 0.388763* 0.447774 0.594927*
* denotes the maximum value
As shown in Table 6, all metrics demonstrate at least moderate correlations ($\rho>0.3$) with human scores on the training set, while exhibiting robust generalization to the test set. In particular, the clue-class and clue metrics achieve strong correlations ($\rho>0.5$). The p-values for all correlations are substantially below 0.05, indicating that the observed associations are statistically significant and reliable. The resulting optimal weight configuration is: $w_{e}=0.810$, $w_{r}=0.190$, $w_{n}=0.787$, $w_{a}=0.104$, $w_{m}=0.109$, $w_{at}=0.594$, $w_{an}=0.406$, $w_{mn}=0.730$, $w_{mt}=0.153$, $w_{mp}=0.117$, $w_{pt}=0.050$, $w_{pn}=0.950$, $w_{rt}=0.156$, $w_{rm}=0.220$, $w_{rn}=0.624$. This configuration will serve as the baseline for subsequent empirical evaluations.
4.4. Metrics Performance Evaluation
In the performance evaluation of our proposed metrics, we selected BLEU (Papineni et al., 2002), CodeBERTScore (Zhou et al., 2023), and NINHS (Al-Khiaty and Ahmed, 2014) as baseline methods for comparison with CLUE. These baselines represent text matching, embedding-based, and domain-specific evaluation approaches, respectively. To assess their efficacy, we computed correlation coefficients between each automatic metric and human ratings derived from the OODEval-Human dataset. A higher correlation coefficient signifies stronger alignment between the automated metric and human preferences, thereby indicating superior metric quality (Evtikhiev et al., 2023; Zhou et al., 2023; Xu et al., 2024). Table 7 reports the Pearson and Spearman correlation coefficients for the baselines, CLUE, and its sub-metrics in relation to human ratings, including their statistical significance levels.
The experimental results reveal two key insights: (1) CLUE substantially surpasses BLEU, CodeBERTScore, and NINHS in both Pearson and Spearman correlations with human ratings, with achieving high statistical significance. In particular, CLUE attains a Pearson correlation coefficient exceeding 0.5, which denotes a strong correlation with human judgments and bolsters interpretability (Cohen, 1992). (2) Relative to its constituent sub-metrics (i.e., clue-class, clue-attribute, clue-method, and clue-relation), the overarching clue metric exhibits superior correlations. This finding underscores that a holistic evaluation encompassing class names, attributes, methods, and relationships aligns more closely with human assessments and offers greater interpretability than evaluations focused on isolated elements.
@@ 4.5. Ablation Study on the Underlying Semantic Model
Since CLUE incorporates a modular semantic model for computing string similarities at the leaf nodes, it is essential to investigate the impact of varying underlying semantic models on its overall performance. To this end, we conducted an ablation study evaluating CLUE across multiple semantic models. Table 8 presents the Pearson correlation coefficients achieved using four representative models, namely MiniLM (Wang et al., 2020), BGE (Chen et al., 2023a), BERT (Devlin et al., 2019), and CodeBERT (Feng et al., 2020) for the CLUE metrics. The results indicate that the choice of underlying semantic model exerts a relatively minor influence on clue-attribute and clue-relation, but a more pronounced effect on clue-method and, particularly, clue-class. Specifically, the range of variation (i.e., the difference between the maximum and minimum correlations) for clue-attribute and clue-relation does not exceed 0.05; for clue-method, it is approximately 0.07; and for clue-class, it reaches about 0.186. Overall, the CLUE metrics grounded in CodeBERT deliver the highest performance across the board. Although the clue-relation and clue-attribute sub-metrics under CodeBERT are marginally inferior to those using BGE, the discrepancies are negligible, with gaps not surpassing 0.025. Consequently, owing to its superior holistic efficacy, we select CodeBERT as the underlying semantic model for all subsequent empirical investigations.
@@ 5. Empirical Study
What are the prevalent failure modes of LLMs on OOD tasks?
@@
Model CL Local? OS? CL? IFT? RLM? PT TDCD
llama3.1-8b-it 128K 2024-07 2023-12
llama3.1-70b-it 128K 2024-07 2023-12
llama2-7b 4K 2023-07 2022-09
llama2-13b 4K 2023-07 2022-09
codellama-7b-it 16K 2023-08 2022-09
codellama-13b-it 16K 2023-08 2022-09
codellama-34b-it 16K 2023-08 2022-09
codellama-7b 16K 2023-08 2022-09
codellama-13b 16K 2023-08 2022-09
codellama-34b 16K 2023-08 2022-09
starcoder2-3b 16K 2024-03 2023-09
starcoder2-7b 16K 2024-03 2023-09
starcoder2-15b 16K 2024-03 2023-09
gemma3-4b-it 128K 2025-03 2024-08
gemma3-12b-it 128K 2025-03 2024-08
gemma3-27b-it 128K 2025-03 2024-08
gemma3-4b-pt 128K 2025-03 2024-08
gemma3-12b-pt 128K 2025-03 2024-08
gemma3-27b-pt 128K 2025-03 2024-08
qwen3-8B 32K 2025-04 2024-11
qwen3-14b 32K 2025-04 2024-11
qwen3-32b 32K 2025-04 2024-11
qwen3-coder-30b 256K 2025-07 2024-11
deepseek-coder-33b 16K 2023-11 2023-03
deepseek-r1 32K 2025-01 2024-07
deepseek-v3 128K 2024-07 2024-07
gpt-3.5-turbo 4K - 2024-03 2022-01
gpt-4o-mini 128K - 2024-07 2023-10
gpt-4o 128K - 2024-05 2023-10
  • CL=Context Length; OS=Open Source; CLM=Code LLM; IFT=Instruct Fine-Tuned; RLM=Reasnoing LLM; PT=Publish Time; TDCD=Training Data Cutoff Date;